The pressure inside a hydrogen-filled container was 2.10 atm

atm at 21 

C

∘C. What would the pressure be if the container was heated to 89 

C

∘C ?


1
Expert's answer
2021-08-22T07:30:28-0400

"P_1=2.1atm, P_2=?"

"T_1=273+21=294K"

"T_2=273+89=362K"

"P_2=\frac{P_1\times T_2}{T_1}=\frac{2.1\times 362}{294}=2.586atm"